0.276 g of Na2CO3.xH2O was weighed out accurately and dissolved in water. Titration with 0.050 mol cm-3 sulphuric acid required 20.00 cm3 for neutralisation. Calculate the value of x.


1
Expert's answer
2020-11-04T14:18:17-0500

Na2CO3 + H2SO4 → Na2SO4 + H2O + CO2

C(H2SO4)= 0.05 mol/L

V(H2SO4) = 20 ml = 0.02 L

n(H2SO4) "= 0.05 \times 0.02 = 0.001 \;mol"

n(H2SO4) = n(Na2CO3) = 0.001 mol

M(Na2CO3) = 106 g/mol

m(Na2CO3) "= n \times M = 0.001 \times 106 = 0.106 \;g"

∆m = 0.276 – 0.106 = 0.17 g

M(H2O) = 18 g/mol

n = m/M

n(H2O) "= \frac{0.17}{18} = 0.0094 \;mol"

"x = \frac{n(H_2O)}{n(Na_2CO_3)} = \frac{0.0094}{0.001} ≈ 10"

Answer: 10.